# CEval: A Benchmark for Evaluating Counterfactual Text Generation

### v1

@@ 3.2 Datasets
To evaluate against human-generated counterfactuals, we selected datasets with contrast sets generated by crowd-sourcing Kaushik et al. (2020) or human experts Gardner et al. (2020). We chose two common benchmark datasets capturing two different NLP tasks: sentiment analysis IMDB Maas et al. (2011) and natural language inference (NLI) on SNLI Bowman et al. (2015). Further datasets with pre-trained classifiers can be easily added to the benchmark.
IMDB contains movie reviews in a diverse range of language styles, sentiments, and topics ensuring that models trained on this dataset are exposed to various linguistic patterns and expressions. Following (Maynez et al., 2023), we select the 488 instances annotated by humans. The dataset is evenly balanced (243 negative and 245 positive instances). We use a pre-trained BERT model [^3] https://huggingface.co/textattack/bert-base-uncased-imdb from TextAttack Morris et al. (2020) as target classifier with an accuracy of 89%. The counterfactual text generation task is to edit the review instances, whether positive or negative, to change the prediction of the classifier on that review with minimal edits. IMDB contains both, contrast examples generated by human experts Gardner et al. (2020) and by crowdworkers Kaushik et al. (2020).
@@
SNLI Bowman et al. (2015) consists of pairs of sentences with a label indicating whether they entail, contradict, or are neutral to each other. NLI requires models to understand the semantic relationships between sentences, which is important for various downstream NLP applications Camburu et al. (2018). Again, we a pre-trained BERT model [^4] https://huggingface.co/textattack/bert-base-uncased-snli from TextAttack Morris et al. (2020) as target classifier for the NLI task with an accuracy of 90%. The counterfactual text generation task is to modify either the premise or hypothesis in the sentence pairs to flip the classifier’s label.
4 Method Selection
In this section, we review state-of-the-art methods for generating counterfactuals, carefully selected from a diverse set of methods found in the literature. To narrow down our comparison in the benchmark, we establish specific properties for method selection. Our criteria aim to make the evaluation comprehensive with minimal effort. We consider not only methods explicitly designed for counterfactual generation but also those with a shared goal of producing high-quality text with respect to a target label. Consequently, we incorporate approaches from the domain of adversarial attacks, expanding the scope beyond traditional counterfactual and contrastive generation methods.
To select methods for the benchmark, we establish the following criteria:
Type of Methods: We categorize methods into Masking and Filling, Conditional Distribution, or Counterfactual Generation with Language Models as mentioned in Section 2. Each category should be covered by a representative method.
Classifier Accessibility: We consider whether the method has access the classifier during the counterfactual generation process.
Reproducibility: We check whether the method provides reproducible source code, ensuring transparency and replicability.
Problem Agnostic: We evaluate whether the method is versatile enough to encompass various types of classification problems.
After filtering based on these criteria (detailed in the Appendix, Table 5), we determine the final set of methods as follows:
@@ 4 Method Selection
MICE Ross et al. (2021) is a contrastive explanation generation method. The idea is, given a predictor and a dataset, MICE trains an editor to generate tokens to fill masked tokens, such as the final text flip the original label. To determine which tokens and how many tokens to mask, the method uses the list of tokens with the highest gradients contributing to the predictions. It employs binary search to find the minimum number of tokens that need to be masked. This method requires access to the classifier to verify the label internally, representing a counterfactual generation method.
@@
LLAMA-2 MICE GBDA CREST Crowd Expert LLAMA-2 MICE GBDA CREST Crowd
CF Metrics Flip Rate $\uparrow$ 0.66 0.99 0.97 0.71 0.81 0.84 0.34 0.85 0.94 0.39 0.75
$\Delta$Probability $\uparrow$ 0.65 0.91 0.96 0.70 0.80 0.83 0.27 0.65 0.86 0.10 0.64
Perplexity $\downarrow$ 41.3 62.1 84.7 44.0 55.9 47.2 54.3 160 143 60.9 72.1
Distance $\downarrow$ 0.23 0.20 0.25 0.33 0.16 0.14 0.58 0.47 0.47 0.32 0.41
Diversity $\uparrow$ 0.24 0.25 0.26 0.43 0.24 0.25 - - - - -
Text Quality Grammar $\uparrow$ 4.80 4.35 3.47 3.89 4.71 4.79 4.83 4.63 3.80 4.21 4.79
Cohesiveness $\uparrow$ 4.63 4.30 3.61 3.53 4.71 4.62 3.67 3.30 2.08 2.83 3.63
Likability $\uparrow$ 2.82 3.12 3.07 2.65 3.12 2.96 3.81 3.31 2.25 3.19 3.52
Fluency $\uparrow$ 3.83 3.58 3.36 3.20 3.81 3.73 4.71 4.60 3.90 4.20 4.67
Average $\uparrow$ 4.02 3.84 3.38 3.32 4.09 4.02 4.25 3.96 3.01 3.61 4.15
Table 2: Results with counterfactual (CF) and text qualily metrics on IMDB and SNLI [^6] We omit diversity, as SNLI has no human expert labels.. Average denotes mean of text quality metrics, each scored on a scale 1-5 following Chiang and Lee (2023)).
GBDA Guo et al. (2021) is a gradient-based adversarial attack, a white-box attack that leverages a novel adversarial distribution formulation, enabling end-to-end optimization of the adversarial loss and fluency constraints with gradient descent. This method requires the use of a model as a classifier, which can be the same as the pretrained model or different, depending on whether access to the original classifier is needed. This method represents the adversarial attack domain.
CREST Treviso et al. (2023) has a similar approach to MICE as they train a rationalizer using SPECTRA Guerreiro and Martins (2021) to find the optimal masks instead of using the highest gradient tokens. Then, they apply filling with T5 following MICE. Given the popularity of Mask and Filling type, we choose another method for a more comprehensive comparison.
LLAMA-2 Touvron et al. (2023): Large Language Models have shown impressive performance on many tasks with only simple prompts Srivastava et al. (2023). Therefore, in this study, we use LLAMA-2 with simple one-shot learning as a baseline for comparing other methods. This choice is made in contrast to other works that used closed-source LLMs.
@@ 5 Results
We evaluate the performance of all compared methods (LLAMA-2, MICE, GBDA, CREST) against human crowd-sourced and human expert generations. It is important to note that MICE and GBDA have access to the prediction model during generation, while CREST leverages a pre-trained T5 model and transfers its findings to the target BERT model. In contrast, LLAMA-2 and both human evaluation groups (crowd and expert) generate counterfactual examples solely based on the provided text and prompt.
Method Text Predicted label
Original If you haven’t seen this, it’s terrible. It is pure trash. I saw this about 17 years ago, and I’m still screwed up from it. Negative (99.94%)
LLAMA-2 If you haven’t seen this, it’s terrible a masterpiece. It is pure trash brilliance. I saw this about 17 years ago, and I’m still in screwed up awe from it. Positive (99.95%)
MICE If you haven’t seen this, it’s terrible pretty. It is pure trash genius. I saw this about 17 years ago, and I’m still screwed up from it. Positive (99.92%)
GBDA If you haven’t seen this, it’s terrible complicated. It is pure trash the magic. I saw it about 17 30 years ago, and I’m still screwed reeling up from it. Positive (99.65%)
CREST If you haven’t seen this movie, it’s terrible definitely worth seeing. It is pure trash ’s great. I saw it about 17 years ago, and I’m still screwed up from it. Positive (99.89%)
Expert If you haven’t seen this, it’s terrible incredible. It is pure trash gold. I saw this about 17 years ago, and I’m still screwed pumped up from it. Positive (99.89%)
Crowd If you haven’t seen this, it’s terrible incredible. It is pure trash gold. I saw this about 17 years ago, and I’m still screwed up hype about it. Positive (99.52%)
Table 3: Example where all methods successfully flipped the label
We start with an illustrative example: Table 3 shows the shortest example in the IMDB dataset where all methods, including human edits, manage to change the predicted label of the original sentence. This example provides a sneak preview on the methods’ varying properties and issues that the quantitative results reveal. For this simple instance, all methods and human groups agree on replacing negative words like terrible and trash with positive words, even though they differ in their choice of positive words. GDBA is the only exception, its replacements do not always convey a positive sentiment, which reduces text quality. Similarly, MICE and CREST fail to detect the negative phrase screwed up, which renders the text less cohesive and fluent then the text generated by LLAMA-2 and humans, who adapt this negative phrase as well. Besides correctly identifying important words, GDBA also replaces irrelevant words like 17 30, resulting in a larger edit distance. For a more complex example with higher variation of edits and generated text, see Table 6 in the Appendix.
@@
The results in Table 6 show that no method outperforms all the others, even when we consider only a single dataset. Expectably, methods with access to the target classifier (MICE & GDBA) are most successful in flipping the label, surpassing even both human groups. However, this success comes at the cost of lower text quality and high perplexity: MICE & GDBA generate “unnatural” text. The fact that humans are less successful in flipping the label indicates limitations within the target classifier, as we would expect a perfect (100%) flip rate for human generation, due to its “gold standard” status. Such potential issues are consistent with prior studies Kaushik et al. (2020); Gardner et al. (2020).
Models and Humans are similarly diverse.
Except for CREST, all methods obtain similar (low) diversity values that are also similar to human diversity values. We attribute these low values to the fact that minimal changes limit the amount of variation. Consequently, CREST has a larger distance to the original text. These observations indicate that if a diverse set of generated text is required, this set should be constructed from the output of different methods.
Counterfactual methods suffer reduced text quality.
In terms of text quality metrics, text generation by humans and LLAMA-2 consistently outperforms other automated methods on most metrics across both datasets. This discrepancy indicates that a focus on counterfactual metrics comes at the cost of reduced text quality. While in particular methods with access to the target classifier (MICE & GDBA) excel at flip rate, they lack in grammar and cohesiveness. Interestingly, LLMs as evaluation proxy seem to prefer the output of another LLM over human generated text. In particular on the SNLI dataset, our evaluation proxy Mistral-7B-instruct consistently prefers LLAMA-2 over human generated text across all text quality metrics. Similarly, GPT-2 as model for measuring perplexity favors LLAMA-2, both on IMDB and SNLI.
@@
Interestingly, MICE has the highest flip rate (FR), but not the largest change in target label probability change ($\Delta$P) on the IMDB dataset. We observe a similar pattern when comparing LLAMA-2 and CREST on the SNLI dataset. CREST has a higher FR, despite LLAMA-2 inducing a larger $\Delta$P.
Figure 1: Histogram of target label probabilities of all methods on the IMDB dataset, including original text and human groups.
The histogram in Figure 1 shows that only MICE generates a significant amount of instances that are close to the decision boundary. All others, including human groups and the original text, exhibit a bimodal pattern with narrow peaks at the two extremes. While the imperfect FR of human groups indicated limitations in the target classifier, the histogram pattern may indicate the source of those limitations: This pattern points to a poorly calibrated, overconfident target classifier, a common issue in today’s deep learning architectures Guo et al. (2017).
LLAMA-2 MICE GBDA CREST Crowd
E N C E N C E N C E N C E N C
Likability 4.29 4.06 2.65 4.18 3.41 2.58 2.90 2.39 1.83 3.56 2.31 1.87 3.79 3.72 3.26
We plan to investigate the impact of a calibrated classifier in future work.
@@ E N C E N C E N C E N C E N C
MICE has the lowest distance among automated methods [^7] In Table 6 we report distance only for true counterfactuals. on the IMDB dataset, but still falls short of human performance. Except for CREST, which has the highest distance on IMDB and lowest on SNLI, distances of all other methods are in a similar range within the same dataset. Despite similar distance ranges, edits might be qualitatively different. To investigate the similarity of edits by different methods, we calculated the average pairwise distance between all generated examples on the IMDB dataset, regardless of label flip success. The results are visualized in Figure 2.
@@
Human crowd-sourced and expert edits are highly similar, indicating substantial overlap in their modifications. LLAMA-2 generated text is closest to human edits, which makes it the most promising candidate to serve as proxy for human-generated counterfactuals. Since LLAMA-2’s edits are so close to human edits, we also attribute part of the flip rate failure to the target classifier, rather than to LLAMA-2 (cf. discussion of expected perfect human flip rate above). GBDA and CREST have the largest distance to all other methods (including the original text) and to each other, i.e., their edits are largely distinct. This substantial difference in generated texts suggests that robustness analyses of the target classifier should always be conducted with multiple methods.
LLMs are a valid evaluation proxy for text quality evaluation.
Previous studies Chiang and Lee (2023); Liu et al. (2023) showed that LLMs can be an alternative to human evaluation. In our results, we observe a tendency of LLMs to prefer the output of other LLMs, both in terms of text quality and perplexity, albeit differences are subtle. To validate Mistral-7B-instruct as text quality evaluation proxy, we compare textual quality metrics of SNLI across two labels: contradiction and entailment. We hypothesize that entailment pairs have higher cohesiveness, fluency, and likability compared to contradicting pairs, as entailment implies a relationship between the two sentences. The results in Table 4 show that entailment pairs score significantly higher in text quality than contradicting pairs, especially in cohesiveness and fluency. We observe consistently higher scores across all methods and human generated text. These findings confirm our hypothesis and we see the same pattern in neutral pairs. As expected, differences in neutral pairs are less pronounced with average values from 2.96 (CREST) to 4.50 (LLAMA-2). [^8] All experiments were conducted with a single NVIDIA A6000 GPU. A full benchmark iteration from counterfactual text generation to evaluation including all methods took approximately 10 days.

### v2

@@ 3.2 Datasets and Classifiers
We chose two benchmark datasets for different NLP tasks: sentiment analysis on IMDB Maas et al. (2011) and natural language inference (NLI) on SNLI Bowman et al. (2015). For both datasets, human-generated counterfactuals from crowdsourcing Kaushik et al. (2020) are available and for IMDB also from experts Gardner et al. (2020). Additional datasets with pre-trained classifiers can be added to the benchmark.
IMDB contains diverse movie reviews from the IMDB website, along with corresponding sentiment labels (positive or negative) for each review. We selected the 488 instances with human-generated counterfactuals, balanced between 243 negative and 245 positive reviews Maynez et al. (2023). Using a pre-trained BERT model [^3] https://huggingface.co/textattack/bert-base-uncased-imdb from TextAttack Morris et al. (2020) with 89% accuracy, the counterfactual task is to minimally edit reviews to alter the classifier’s prediction.
@@
SNLI Bowman et al. (2015) consists of sentence pairs labeled as entailment, contradiction, or neutral, requiring models to understand semantic relationships. Using a pre-trained BERT model [^4] https://huggingface.co/textattack/bert-base-uncased-snli from TextAttack Morris et al. (2020) with 90% accuracy, the counterfactual generation methods have to modify the premise or the hypothesis to change the classifier’s label.
4 Counterfactual Methods Selection
In this section, we briefly describe the counterfactual generation methods we evaluate with our benchmark. We selected at least one representative for each of the categories Masking and Filling (MF), Conditional Distribution (CD) and Large Language Models (LLMs) (cf. Section 2) based on the following criteria:
  • The authors provide reproducible source code.
  • The method is problem agnostic and can be applied to multiple text classification tasks.
  • The method has access to the underlying text classifier.
We used the criteria reproducible code and problem agnostic as hard filters and access to the target classifier as soft filter. A problem agnostic method is versatile enough to generate counterfactuals for various types of classification problems (whereas methods like Polyjuice Wu et al. (2021) or Tailor Ross et al. (2022) require control codes, which limits their flexibility). Methods without access to the target classifier are at disadvantage, as they have no information about the internals of the target classifier. Hence, wherever available, we opted for a method with access to the target classifier. The selection based on these criteria (cf. details in Appendix, Table 4) resulted in MICE, GDBA, CREST and LLAMA-2 as representative counterfactual generation methods. We briefly describe them in the following.
@@ 4 Counterfactual Methods Selection
MICE Ross et al. (2021) is a contrastive explanation generation method. It trains an editor to fill masked tokens in a text so that the final text changes the original label. The tokens to be masked are chosen based on the highest gradients contributing to the predictions, and binary search is used to find the minimum number of tokens to mask. This method requires access to the classifier to verify the label internally, representing a counterfactual generation method.
@@
LLAMA-2 MICE GBDA CREST Expert Crowd LLAMA-2 MICE GBDA CREST Crowd
CF Metrics Flip Rate $\uparrow$ 0.7 1.0 0.97 0.71 0.81 0.85 0.39 0.85 0.94 0.39 0.75
$\Delta$Probability $\uparrow$ 0.69 0.91 0.96 0.70 0.80 0.84 0.33 0.65 0.86 0.10 0.64
Perplexity $\downarrow$ 41.3 62.1 84.1 44.7 56.2 52.4 57.0 160 143 60.9 72.1
Distance $\downarrow$ 73.9 38.5 46.1 70.5 29.3 25.0 6.15 5.64 4.85 3.53 4.06
Diversity $\uparrow$ 61.6 48.4 47.6 86.6 38.7 38.7 - - - - -
Text Quality Grammar $\uparrow$ 3.18 2.71 2.16 2.18 2.90 2.92 3.68 3.33 2.29 2.71 3.58
Cohesiveness $\uparrow$ 3.12 2.81 2.38 2.27 2.99 2.95 3.61 3.31 2.03 2.74 3.60
Fluency $\uparrow$ 3.13 2.79 2.37 2.33 2.99 2.92 3.59 3.33 2.17 2.70 3.56
Average $\uparrow$ 3.14 2.77 2.30 2.27 2.96 2.93 3.63 3.33 2.16 2.72 3.58
Table 2: Results with counterfactual (CF) and text quality metrics on IMDB and SNLI. Average denotes average of text quality metrics, each scored on a scale 1-5 following Chiang and Lee (2023). We calculate diversity of the human groups by comparing expert with crowd counterfactuals and omit diversity on SNLI as it only has a single human counterfactual per instance (no expert annotations).
GBDA Guo et al. (2021) is a gradient-based adversarial attack that uses a novel adversarial distribution for end-to-end optimization of adversarial loss and fluency constraints via gradient descent. Similar to MICE, this approach needs access to the classifier for internal label verification. This method represents the adversarial attack domain.
CREST Treviso et al. (2023) follows a similar approach as MICE in first masking tokens that should be changed. Instead of using the highest gradient tokens to find the masks, the authors train a rationalizer using SPECTRA Guerreiro and Martins (2021). Then, they fill the blanks with T5 same as MICE. Given the popularity of the Mask and Filling type, we chose this method for a more comprehensive comparison.
LLAMA-2 Touvron et al. (2023): Large Language Models have shown good performance on many tasks with only simple prompts Srivastava et al. (2023). Therefore, in this study, we use LLAMA-2 with simple one-shot learning as a baseline that is not specifically designed for counterfactual generation, but has strong language generation capabilities. The choice for LLAMA-2 as an open-source model is made in contrast to other studies that used closed-source LLMs.
The hyperparameters of each selected method can significantly impact the results, particularly for MICE Ross et al. (2021) and CREST Treviso et al. (2023). The percentage of masked tokens in both methods, representing the upper bound of changed tokens, directly influences the token distance and indirectly affects the flip rate: a lower percentage allows fewer tokens to change, resulting in a smaller distance but potentially a lower flip rate. In our experiments, we maintain the hyperparameters as specified in the original papers of each method. In case of LLAMA-2, the temperature of LLMs affects word sampling: lower temperatures yield more deterministic results, while higher temperatures enhance creativity. For the comparison with other methods, we use a temperature of 1.0 and analyze the impact of varying temperatures at the end of the next section.
@@ 5 Results
We evaluate all counterfactual generation methods against human crowd-sourced and human expert generations. Note that MICE and GBDA have access to the prediction model during generation, while CREST employs a pre-trained T5 model for internal label verification and transfers its prediction to the target BERT model. In contrast, LLAMA-2 and both human evaluation groups (crowd and expert) generate counterfactual examples solely based on the provided text and prompt.
We start with an example to illustrate the methods’ varying characteristics before discussing our observations from the quantitative results. Fig. 1 shows the shortest example in the IMDB dataset where all methods, including human edits, change the label of the original sentence on the generated counterfactual. For this simple instance, all methods and human groups agree on replacing negative words like terrible and trash with positive words, even though they differ in their choice of positive words. GDBA is the only exception, its replacements do not always convey a positive sentiment, which reduces text quality. Similarly, MICE and CREST fail to detect the negative phrase screwed up, which renders the text less cohesive and fluent than the text generated by LLAMA-2 and humans, who adapt this negative phrase as well. Besides correctly identifying important words, GDBA also replaces irrelevant words like 17 30, resulting in a larger edit distance. For a more complex example with higher variation of edits and generated text, see Table 9 in the Appendix.
@@
Table 2 shows that no single method consistently outperforms the others, even on a single dataset. Methods with access to the target classifier, such as MICE and GDBA, excel at flipping the label but generate “unnatural” text with lower quality and higher perplexity due to poor grammar and low cohesiveness. In contrast, humans and LLAMA-2 consistently produce higher quality text across most metrics on both datasets. The lower success rate of humans in flipping the label suggests limitations in the target classifier, as perfect flip rates would be expected for human-generated text, the “gold standard.” Such potential issues are consistent with prior studies Kaushik et al. (2020); Gardner et al. (2020). Additionally, LLMs used as evaluation proxies, such as ChatGPT and GPT-2 (which measures perplexity), prefer LLAMA-2’s output over human-generated text on both the SNLI and IMDB datasets. This preference is observed across different evaluator temperatures, as shown in Table 3, suggesting an interesting direction for further research into bias of LLMs as evaluators.
Diversity and distance are correlated.
On the IMDB dataset, CREST and LLAMA-2 exhibit the highest diversity but also the highest distance. In contrast, human-generated changes (crowd and expert) are minimal and the least diverse. The Pearson correlation between diversity and distance is 0.93, indicating a very strong correlation between these two metrics. This strong correlation is likely due to minimal changes limiting the amount of variation.
@@
Interestingly, MICE has the highest flip rate (FR), but not the largest change in target label probability change ($\Delta$P) on the IMDB dataset. We observe a similar pattern when comparing LLAMA-2 and CREST on the SNLI dataset. CREST has an equal FR, despite LLAMA-2 inducing a larger $\Delta$P.
Figure 2: Distribution of target label probabilities of all methods on the IMDB dataset, including original text and human groups.
A high FR combined with a low $\Delta$P suggests that the counterfactuals generated by the method are close to the decision boundary of the target classifier. Fig. 2 shows that only MICE generates a noticeable amount of instances that are close to the decision boundary ($P(y^{\prime})=0.5$). All others, including human groups, exhibit a bimodal pattern with narrow peaks at the two extremes. While the imperfect FR of human groups suggests limitations in the target classifier, the distribution pattern may indicate the source of those limitations: This pattern points to a poorly calibrated, overconfident target classifier, a common issue in today’s deep learning architectures Guo et al. (2017).
@@
Among automatically generated methods, MICE’s counterfactuals are closest to the original texts [^5] In Table 2 we report distance only for true counterfactuals. on the IMDB dataset, but still edit more tokens than humans (expert and crowd). The distance scores of CREST and LLAMA-2 are similar, as are those for MICE and GBDA, and for expert and crowd edits on the IMDB dataset. However, similar edit distances do not imply that these methods make the same edits. To investigate the similarity of edits by different methods, we calculated the average pairwise distance between all generated examples on the IMDB dataset, regardless of label flip success. The results are visualized in Fig. 3.
@@
Crowd and expert edits are highly similar, indicating substantial overlap in their modifications. MICE generated text is closest to human edits, which makes it the most promising candidate to serve as proxy for human-generated counterfactuals. GBDA and CREST have the largest distance to all other methods (including the original text) and to each other, i.e., their edits are largely distinct. This substantial difference in generated texts suggests that robustness analyses of the target classifier should always be conducted with multiple methods.
Grammar Cohesiveness Fluency
GPT Mistral GPT Mistral GPT Mistral
0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0 0.2 1.0
Expert 2.90 2.94 4.81 4.74 2.99 2.99 4.74 4.66 2.99 2.99 3.91 3.91
Crowd 2.92 2.89 4.88 4.79 2.95 2.98 4.78 4.68 2.92 2.94 3.83 3.81
Crest 2.18 2.15 4.05 3.96 2.27 2.30 3.95 3.91 2.33 2.37 3.36 3.34
GBDA 2.16 2.18 3.92 3.82 2.38 2.40 4.00 3.89 2.37 2.35 3.44 3.46
Mice 2.71 2.73 4.55 4.44 2.81 2.82 4.40 4.35 2.79 2.81 3.77 3.75
LLAMA-2 3.18 3.19 4.90 4.86 3.12 3.11 4.83 4.74 3.13 3.12 4.00 3.96
Table 3: Comparison of text quality evaluation using Mistral and ChatGPT (GPT-3.5 Turbo) with different temperatures (0.2 and 1.0) on IMDB dataset.
Temperature affects counterfactual generation diversity
We compare LLAMA-2’s temperature setting of 1.0 in Table 2 with additional values of 0.2 and 0.6 for counterfactual generation and observe that the diversity score of LLAMA-2 varies significantly with temperature changes: the lower the temperature, the lower the diversity. For a temperature of 0.2, diversity score is 28.3 and for temperature 0.6, diversity score is 44.4 (details in Appendix, Table 6). This finding aligns with the expectation that higher temperatures, which increase token sampling flexibility, enhance the diversity of generated text. In contrast, other metrics remain largely unchanged or show minor variations. For instance, average text quality is 3.15 at both temperatures of 0.6 and 0.2 on IMDB dataset.
6 Comparison of LLMs for Text Quality Evaluation
Evaluating text quality with ChatGPT has been shown to be effective Huang et al. (2023); Gilardi et al. (2023). However, such evaluations come at high costs, limited control and customization constraints, and lack transparency. Therefore, we investigate an open-source LLM, Mistral-7B Jiang et al. (2023) as an evaluation proxy.
Mistral-7B is a valid alternative to ChatGPT
To validate Mistral’s evaluation capability, we use Mistral to evaluate the counterfactuals generated by all methods and compare the assessment scores with those from ChatGPT. Specifically, we compare the average scores, the Pearson correlation on the scores of each instance, and the Spearman correlation of the ranking of each method on all text quality metrics on both datasets and two temperature settings of 0.2 and 1.0. Table 3 shows that Mistral-7B generally assigns higher scores than ChatGPT across all text quality metrics, though their scores are correlated. The Pearson correlation on the scores of each instance from the two models ranges from moderate to strong, with coefficients from 0.4 to 0.7, regardless of temperature settings (details in Appendix, Fig. 4). This implies that a text with high scores from Mistral is likely to receive high scores from ChatGPT as well. Furthermore, Spearman’s rank correlation coefficients on the scores between the two models range from 0.89 to 1.0 , indicating a very strong correlation and partly even exactly identical rankings (details in Appendix Table 5).
To further validate Mistral-7B-instruct as a text quality evaluation proxy, we analyzed textual quality metrics on SNLI across two labels: contradiction and entailment. We hypothesized that entailment pairs exhibit higher cohesiveness and fluency than contradiction pairs, as entailment implies a logical relationship between the sentences. Our evaluation confirms that entailment pairs score significantly higher in text quality, particularly in cohesiveness and fluency, across all methods and human-generated texts. Detailed results are provided in Appendix, Table 7.
Given the moderate to strong correlation with ChatGPT scores, very strong correlation in rankings and the validation of textual quality on the SNLI dataset, Mistral-7B is a viable alternative for comparative counterfactual method evaluation.
Text quality evaluation is robust to temperature variations
Since temperature influences the performance of LLMs during inference Wang et al. (2023a), we evaluate its impact on their evaluation capabilities. Our study finds that text quality evaluation results are robust to temperature changes for both Mistral-7B and ChatGPT. We find a very strong correlation (Pearsons $\rho$ > 0.8) between evaluation scores for different temperatures of the same model (Appendix Figures 4 and 5). Furthermore, the absolute scores remain similar across temperatures, as shown in Table 3.